OpenStudy (anonymous):

findd2y/dx2 in terms of 3x^2-2y^2=5

OpenStudy (yttrium):

Implicit differentiation twice. 1st: 6x - 4yy' = 0 6x = 4yy' y' = 3x/2y y'' = ((2y(3)-(3x(2y'))/4y^2 y'' = (6y-6xy')/4y^2 y'' = 3(y - x (3x/2y) )/2y^2 y'' = 3(2y^2 - -3x^2)/2y^3